The greater of the two consecutive even numbers is 6 less than twice the smaller.
Answer by haileytucki(390)

okay, you need 2 consecutive integers (x+2)=(2x-6) would be your initial set up.
You can plu in any even numbers to complete this problem, so long as they are even. If the problem were to be worked out;
(x+2)=(2x-6)
Remove the parentheses around the expression x+2.
x+2=(2x-6)
Remove the parentheses around the expression 2x-6.
x+2=2x-6
Since 2x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 2x from both sides.
x+2-2x=-6
Since x and -2x are like terms, add -2x to x to get -x.
-x+2=-6
Since 2 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2 from both sides.
-x=-2-6
Subtract 6 from -2 to get -8.
-x=-8
Multiply each term in the equation by -1.
-x*-1=-8*-1
Multiply -x by -1 to get x.
x=-8*-1
Multiply -8 by -1 to get 8.
x=8

Please let me know if you don't follow, or email me at Jennifer_Sadler@stu.southuniversity.edu